1) A+B+C=360 THEN COT A/4 + COT B/4 +COT C/4 = ​2) SINx COSx =1/4 AND 2TANx =4 TANy SIN (x-y) =

1)  A+B+C=360 THEN COT A/4 + COT B/4 +COT C/4 =
​2) SINx COSx =1/4 AND 2TANx =4 TANy SIN (x-y) =

Grade:11

1 Answers

Arun
5 years ago
Dear student
 
Please ask only one question in one thread
1.
when A+B+C=90 or 360 
A+B+C=90
A+B=90-C
cot(A+B)=Cot(90-C)
cotAcotB-1=cotA+cotB/cotC
cotAcotBcotC-cotC=cotA+cotB
cotA+cotB+cotC=cotAcotBcotC
cotA+cotB+cotC=cotAcotBcotC  
cotA/4+cotB/4+cotC/4=cotA/4cotB/4cotC/4